# Tuning LLM Judges Hyperparameters

### v1

@@ Abstract
Evaluating Large Language Models (LLMs) often requires costly human annotations. To address this, LLM-based judges have been proposed, which compare the outputs of two LLMs enabling the ranking of models without human intervention. While several approaches have been proposed, many confounding factors are present between different papers. For instance the model, the prompt and other hyperparameters are typically changed at the same time making apple-to-apple comparisons challenging. In this paper, we propose to systematically analyze and tune hyperparameter of LLM judges. To alleviate the high cost of evaluating a judge, we propose to leverage multi-objective multi-fidelity which allows to find judges that trades accuracy for cost and also reduce significantly the cost of the search. Our method identifies judges that not only outperform existing benchmarks in accuracy and cost-efficiency but also utilize open-weight models, ensuring greater accessibility and reproducibility.
@@ 1 Introduction
Instruction tuned models are difficult to evaluate given that they provide free-form text given arbitrary instructions that may include summarization, code writing or translation. While humans can be used to annotate the quality of the outputs of an LLM given instructions, it becomes quickly expensive and also delays the evaluation and development of instruction-tuned models (Li et al., 2023).
As an alternative, LLM judges have been proposed to provide indicative ranking for instruction-tuned models (Li et al., 2023), but also to select instruction-tuned recipes (et. al., 2024; Lambert et al., 2024). While they can reduce cost significantly, LLM judges have also limitations as they may rely on superficial style aspect such as the length of a response (Dubois et al., 2024) or the order of their input (Zheng et al., 2023).
While some of those issues can be easily fixed, several aspects limit the improvement of LLM judges. The first is that evaluating a LLM judge configuration is expensive, for instance evaluating one model in Alpaca-Eval (Li et al., 2023) costs $\sim$24\$ (Ni et al., 2024). However, to evaluate a judge, one has to generate annotations on many models in order to compute the Spearman correlation with Elo ratings. Given this, many confounding factors are typically present between two different judges approach. For instance, between Alpaca-Eval (Li et al., 2023) and Arena-Hard judges (Li et al., 2024), the judge model, the prompt, the score-type and the set of instructions were changed. This makes it hard to learn which contributions are important.
In this work, we propose to analyze and tune systematically the hyperparameters of a LLM judge, including the LLM model used, the prompt, the inference parameters (such as the temperature) as well as the parsing mechanism used to extract the judge preference. We first analyze the impact of scaling the LLM model or the number of instructions on the judge performance. This highlights that scaling is insufficient to reach good performance and also helps us to identify a cheaper way to evaluate judges than evaluating Spearman correlation with Chatbot Arena with a grid of model annotations. We then show how one can systematically tune the hyperparameters of a judge using a multi-objective (to account for accuracy and cost) and multi-fidelity (to save tuning cost) approach. Finally, we show that the configurations found outperform previous state-of-the-art judges on a range of real-world test-datasets.
@@
  • We study scaling laws of judges showing how much model-size and number of instructions alone impact key metrics used to evaluate judges
  • We propose a way to tune judges at reasonable tuning cost
  • We show the approach is able to identify configuration that outperform previous approaches while relying solely on open-weight models
  • We analyze which prompt strategy and hyperparameters work best for judges highlighting import patterns that may be used by the community to build better judges
@@ 2 Related work
LLM as a judge has been emerging as a way to alleviate large human annotation costs that are required to evaluate instruction tuned models. For instance, Llama3 used an earlier version of itself as a reject sampler to select best completions (et. al., 2024) or more recently Tulu (Lambert et al., 2024) used an LLM judge to annotate preference data to perform instruction tuning. LLM judges have also been used to provide leaderboards such as Alpaca-Eval (Li et al., 2023) or Arena-Hard (Li et al., 2024) which offer cheaper alternative than human-annotated leaderboards such as ChatBot Arena (Chiang et al., 2024).
Zeroshot and finetuned judges.
Two main approaches have been proposed for LLM judges. The first one, referred as zeroshot-judges, prompts a LLM to rate a pair of model completions (Li et al., 2023, 2024) or single completion possibly against a baseline (Zheng et al., 2023). The second one, refered to as finetuned-judges, fine-tunes an existing model on a set of human annotated preferences from pair of outputs (Zhu et al., 2023; Wang et al., 2024). In contrast with the first approach, it requires finetuning a model which may occur an extra-cost, it has also been observed than fine-tuned judges are less robust under distribution shift (Huang et al., 2024).
Zeroshot judges.
Many strategies for zeroshot judges have been proposed. Li et al. (2024) proposed to ask the judge to answer the instruction to perform some form of Chain of Thought (Wei et al., 2022). To avoid the order of outputs to matter in pairwise comparison, previous work proposed to randomize or average the two possible positions (Dubois et al., 2024; Li et al., 2024). To parse the model output, Li et al. (2023) asks the judge a letter to denote the best model, (Li et al., 2024) uses instead a Likert scale (such as A>>B, A>B, A=B to indicate respectively when model A is much better, better or comparable wrt B), and other alternative such as (Cui et al., 2024; Lambert et al., 2024) outputs a score for various criteria such as Instruction-following, honesty, or helpfulness. The underlying LLM models is often changed from one paper to the other in addition to all those dimension which does not allow comparing which strategy inherently work better.
@@
In parallel with their adoption, several limitations of LLM judges have been highlighted. Among them are their dependence to superficial stylistic aspect of an answer, for instance favoring longer answers (Dubois et al., 2024), favoring the first answer when using judges that make pairwise comparison (Li et al., 2024), or more problematically favoring its own output (Panickssery et al., 2024).
While some of those issues have been fixed by averaging the order of the answers (Li et al., 2024) or using a causal model to isolate the length effect from the judge preference (Dubois et al., 2024), the problem of favoring its own answer is more problematic. Indeed, most previous work use judges from close models such as GPT-4 which may bias leaderboard or simply make them unavailable if model stop being available or their price increase.
@@
To the best of our knowledge, tuning the hyperparameters of judges have not been fully investigated. One reason is the associated compute cost with naive approaches. For instance, Alpaca-Eval and Arena-Hard evaluates each judge configuration on a grid of models and instructions causes a significant cost when comparing multiple judges [^1] For instance, to evaluate $4480$ judges configurations as we do in this work, we estimate that the cost would exceed $2$M\$ for either Alpaca-Eval or Arena-Hard, see note B.3 in appendix.. Cost is therefore a strong limitation for the tuning of judges limiting the comparison and tuning of judges to simple aspect suchs as the choice of the underlying LLM model but excluding other key factors such as the prompt strategy, the output format or the temperature.
In this work, we propose to search for judge configurations including the optimal prompt parametrization. Prompt tuning approaches such as (Fernando et al., 2023) are related although they did not consider tuning judges. One main difference is that we focus not only on tuning the prompt but also on other judge hyperparameters such as the model or the temperature. Similarly to our work, Doddapaneni et al. (2024) analyzed the effectiveness of 5 different prompt strategies for judges. In contrast to our work, the approach is focused on introducing a benchmark whereas in our case we focus on tuning and analyzing judges in a search space containing 80 prompting strategy and 4480 judge configurations in total.
We begin by providing background on LLM judges and then explore how much improvement can be obtained through scaling alone. This analysis highlights the importance of human agreement as a more efficient metric for distinguishing between judges. We then show how one can tune judges including their prompt and other hyperparameters such as their model or temperature.
@@ 3.1 Judge
inline]Reuse jury (Verga et al., 2024) paper exposition for pair judges We denote a LLM model as a function $\pi:p\to o$ that produces an output string $o$ given a prompt $p$. A judge compares two models $\pi_{0}$ and $\pi_{1}$ and outputs a score $\Phi(p,\pi_{0}(p),\pi_{1}(p))\in[0,1]$ which is close to zero if $\pi_{0}(p)$ is better than $\pi_{1}(p)$ or close to one otherwise.
A common choice is to use a fix baseline $\pi^{*}$ for one of the model, typically a frontier model which allows to obtain a score for a model to be evaluated: $\Phi(p,\pi(p))=\Phi(p,\pi^{*}(p),\pi(p))$. Given that the order can sometimes influence the judge decision (Li et al., 2023), previous works have proposed to either randomize the baseline position (Li et al., 2023) or compute the judgement given both order and average out the result (Li et al., 2024).
@@
One approach to see how well a judge performs is to check how close are its rankings on a grid of models and instructions compared to ranking obtained through human annotations.
Assume we are considering a finite set of models $\mathcal{M}$ and instructions $\mathcal{P}$. To evaluate a judge, we also assume that we are given a list of golden scores for the models denoted $s^{h}\in\mathbb{R}^{\mathcal{M}}$, for instance the elo-ratings from the Chatbot Arena obtained from human ratings.
We can then compare the golden scores with the scores $s^{\Phi}\in\mathbb{R}^{\mathcal{M}}$ obtained by averaging the judge performance across all instructions, e.g. $s^{\Phi}_{i}=\mathbb{E}_{p\sim\mathcal{P}}\left[\Phi(p,\pi_{i}(p))\right]$.
One possible metric to measure a judge quality is the Spearman correlation with the golden score, e.g.
$\rho(s^{h},s^{\Phi})=\frac{\text{cov}\left[\text{R}(s^{h}),\text{R}(s^{\Phi})\right]}{\sigma(\text{R}(s^{h}))\sigma(\text{R}(s^{\Phi}))}\in[-1,1]$ (1)
where $\text{R}(x)$ and $\sigma(x)$ denotes respectively the rank operator and the standard deviation. Using Spearman correlation can be beneficial as the scales from the golden scores and the scale from the judge annotation may be different. Other metrics which can be used to compare the order of models include Brier score and calibration described (Li et al., 2024).
@@
Another way to evaluate judges consists in measuring their agreement with preferences annotated by humans in a list of pairwise model battles. A battle consists in a prompt, a pair of model outputs, and binary indicating which model was preferred by the human annotator.
@@
$\mathbb{E}_{i\sim N}[\Phi(p,o_{i},o^{\prime}_{i})=\Phi^{\text{h}}(p_{i},o_{i},o^{\prime}_{i}))]$ (3)
inline]Another approach consists in looking at Cohen’s kappa score, for instance done in (Verga et al., 2024) however numbers not comparable as they use QA which is a much easier task.
inline](Verga et al., 2024) also compared with CB Arena, looking at a few prompt variation, no tuning.
@@ 4 Scaling judges
Figure 1: Effect on scaling the LLM judge and the number of instructions on Spearman correlation. In constrast with human-agreement, neither Alpaca-Eval, Arena-Hard or the union of both allow to separate the quality between 32B and 72B models.
Rather than tuning judges a natural question is: can we just scale them up? LLM Judges can be scaled in multiple ways: by increasing the number of parameters of the LLM judge or by using instructions. Here we investigate how well the Spearman correlation and human agreement metrics scale with both dimensions.
@@
In Fig. 1, we show the scaling behavior when increasing the LLM judge size and the number of instructions. We use Qwen2.5 with a default prompt given in the appendix and compute Spearman correlation on a common set of 24 models available on both Alpaca-Eval and Arena-Hard.
As expected, the judge performance improves when scaling the LLM model and when using more instructions as it allows to cover more area of the models to evaluate. Alpaca-Eval contains easier prompts and therefore gives better performance to smaller judges compared to Arena-Hard. In contrast, Arena-Hard contains harder and technical question. This allows to reach better Spearman correlation as the instructions allow to separate more advanced models, provided that the judge base model is strong enough to measure the performance on this more complex set of instructions.
@@
In Fig. 2, we study the effect of scaling the LLM judge and the number of battles this time on human agreement. Here, we use the same prompt and base LLM models as in the previous paragraph but compute human agreement on instructions of LMSys dataset (lin Chiang et al., 2024).
@@
Model-size Spearman corr. Human agreement
0.5B 0.12 $\pm$ 0.188 (0.74\$) 0.36 $\pm$ 0.006 (0.45\$)
1.5B 0.33 $\pm$ 0.125 (0.74\$) 0.37 $\pm$ 0.006 (0.46\$)
3B 0.82 $\pm$ 0.064 (1.24\$) 0.42 $\pm$ 0.006 (0.76\$)
7B 0.84 $\pm$ 0.080 (3.87\$) 0.45 $\pm$ 0.006 (2.38\$)
32B 0.90 $\pm$ 0.051 (10.24\$) 0.48 $\pm$ 0.006 (6.30\$)
72B 0.88 $\pm$ 0.072 (15.45\$) 0.50 $\pm$ 0.006 (9.51\$)
Table 1: Relation with Spearman-correlation with Chatbot Arena and Human-agreement for all sizes of Qwen on a default prompt. We evaluate 400 instructions (times 24 models) for Spearman correlation and 6400 instructions to compute human agreement.
@@
We have two metrics to evaluate the judge quality: the Spearman correlation and the human agreement. As discussed in the previous paragraph, they behave differently as human agreement is an average of an atomic property (does the judge and human agrees on instruction on average) whereas Spearman correlation is a global metric. Both metrics are correlated, we can see for instance that the ranking w.r.t. the base model size used for the judge is mostly consistent for both Spearman correlation and human agreement in Figs. 1 and 2 and for both metrics, we observe that, for this prompt and model family, LLM judges bellow 7B are not able to outperform a simple baseline length which favor the output with the longest answer.
In what follows, we use human agreement as the metric to optimize as it allows to rank judge configurations much more cheaply than Spearman correlation and has a better signal-to-noise ratio. Indeed, it can be seen in Figs. 1 and 2 that judge configurations are separated with fewer instructions when using human-agreement. In Table 1, we sample instructions and report standard deviation for both metrics. It can be seen that while both metrics are correlated, human-agreement allows to differentiate larger models much more closely as the signal to noise ratio of the metric is better. In particular, it allows to statistically distinguish 32B and 72B models whereas those models are tied for Spearman correlation given the number of instructions considered.
inline]Table 1 is an apple to orange comparison. Sure we use less evaluations for human agreement and it is better but ot would be clearer to match the number of evaluations, try to evaluate 200 instr x 24 models and 5000 LMSys which would match the number of evaluations.

### v2

# Tuning LLM Judge Design Decisions for 1/1000 of the Cost

@@ Abstract
Evaluating Large Language Models (LLMs) often requires costly human annotations. To address this, LLM-based judges have been proposed, which compare the outputs of two LLMs enabling the ranking of models without human intervention. While several approaches have been proposed, many confounding factors are present between different papers. For instance the model, the prompt and other hyperparameters are typically changed at the same time making apple-to-apple comparisons challenging. In this paper, we propose to systematically analyze and tune the hyperparameters of LLM judges. To alleviate the high cost of evaluating a judge, we propose to leverage multi-objective multi-fidelity which allows to find judges that trade accuracy for cost and also significantly reduce the cost of the search. Our method identifies judges that not only outperform existing benchmarks in accuracy and cost-efficiency but also utilize open-weight models, ensuring greater accessibility and reproducibility.
@@ 1 Introduction
Instruction tuned models are difficult to evaluate as they provide free-form text given arbitrary instructions that may include summarization (Zhang et al., 2024), code writing (Ni et al., 2023) or translation (Elshin et al., 2024). While humans can annotate the quality of the outputs of an LLM given instructions, this quickly becomes expensive and also delays the evaluation and development of instruction-tuned models (Li et al., 2023).
As an alternative, LLM judges have been proposed to provide an indicative ranking for instruction-tuned models (Li et al., 2023), but also to select instruction-tuning recipes (Grattafiori et al., 2024; Lambert et al., 2024). While they can reduce cost significantly compared to human evals, LLM judges have limitations as they may rely on superficial style aspects, such as the length of a response (Dubois et al., 2024) or the order of their input (Zheng et al., 2023).
While some of these issues can be easily fixed, several aspects limit the improvement of LLM judges. The first is that evaluating a LLM judge configuration is expensive. For instance, evaluating one model in Alpaca-Eval (Li et al., 2023) costs $\sim$24\$ (Ni et al., 2024), and evaluating a judge multiplies this by the large number of model evaluations it has to do in order to compute the correlation with human ratings. Given this, many confounding factors are typically present between different judges approaches. For instance, between Alpaca-Eval (Li et al., 2023) and Arena-Hard (Li et al., 2024), the judge model, the prompt, the score-type and the set of instructions were changed. This makes it hard to learn which contributions are important.
In this work, we propose to analyze and tune systematically the hyperparameters of an LLM judge, including the LLM model used, the prompt, the inference parameters (such as the temperature) as well as the parsing mechanism used to extract the judge preference. We first analyze the impact of scaling the LLM judge model or the number of instructions on the judge performance. This highlights that scaling is insufficient to reach good performance and also helps us to identify a cheaper way to evaluate judges than evaluating Spearman correlation with Chatbot Arena with a grid of model annotations. We then show how to systematically tune the hyperparameters of a judge using a multi-objective (to account for accuracy and cost) and multi-fidelity approach which saves tuning cost by stopping early poor configurations. Finally, we show that the configurations found outperform previous state-of-the-art judges on a range of real-world test datasets.
@@
  • We study scaling laws of judges showing how much model size and number of instructions alone impact key metrics used to evaluate judges
  • We propose a way to tune judges hyperparameters at reasonable tuning cost
  • We show the approach is able to identify configurations that outperform previous approaches while relying solely on open-weight models
  • We analyze which prompt strategy and hyperparameters work best for judges highlighting important patterns that may be used by the community to build better judges
@@ 2 Related work
LLM as a judge has been emerging as a way to alleviate large human annotation costs that are required to evaluate instruction-tuned models. For instance, Llama3 used an earlier version of itself as a reject sampler to select best completions (Grattafiori et al., 2024) or more recently (Lambert et al., 2024) used an LLM judge to annotate preference data to perform instruction tuning. LLM judges have also been used for leaderboards such as Alpaca-Eval (Li et al., 2023) or Arena-Hard (Li et al., 2024) which offer cheaper alternatives than human-annotated leaderboards such as ChatBot Arena (Chiang et al., 2024).
Zero-shot and fine-tuned judges.
Two main approaches have been proposed for LLM judges. The first one, referred to as zero-shot judges, prompts a LLM to rate a pair of model completions (Li et al., 2023, 2024) or a single completion possibly against a baseline (Zheng et al., 2023). The second one fine-tunes an existing model on a set of human-annotated preferences (Zhu et al., 2023; Wang et al., 2024). In contrast with the first approach, it requires fine-tuning a model which may occur at an extra cost, along with lower robustness under distribution shift (Huang et al., 2024).
Zero-shot judges.
Many strategies for zero-shot judges have been proposed. Li et al. (2024) proposed to ask the judge to answer the instruction to perform some form of Chain of Thought (Wei et al., 2022). To avoid the order of outputs to matter in pairwise comparison, previous work proposed to randomize or average the two possible positions (Dubois et al., 2024; Li et al., 2024). To parse the model output, Li et al. (2023) asks the judge a letter to denote the best model, Li et al. (2024) uses instead a Likert scale (such as A>>B, A>B, A=B to indicate respectively when model A is much better, better or comparable wrt B), and another alternative (Cui et al., 2024; Lambert et al., 2024) outputs a score for various criteria such as instruction-following, honesty, or helpfulness. The underlying LLM models often vary between papers, along with multiple other dimensions, making it difficult to determine which strategies inherently perform better.
@@
In parallel with their adoption, several limitations of LLM judges have been highlighted. Among them is their dependence on superficial stylistic aspects of an answer, for instance favoring longer answers (Dubois et al., 2024), the first answer when using judges that make pairwise comparison (Li et al., 2024), or their own outputs (Panickssery et al., 2024).
While some of those issues have been fixed by averaging the order of the answers (Li et al., 2024) or using a causal model to isolate the length effect from the judge preference (Dubois et al., 2024), the challenge of favoring its own answer is more problematic. Indeed, many previous works used judges from close models such as GPT-4 which may bias the leaderboard or render evaluations infeasible if the model becomes unavailable or its cost increases.
@@
To the best of our knowledge, hyperparameter tuning for LLM judges has not been comprehensively explored in a way that accounts for the various factors contributing to a high-performing judge. One reason is the associated compute cost with naive approaches. For instance, Alpaca-Eval and Arena-Hard evaluate each judge configuration across a grid of models and instructions, leading to substantial expenses when comparing multiple judges [^1] We estimate that our approach costs approximately 2K\$ to search through 4,480 judge configurations and that evaluating the same number of judges using Alpaca-Eval or Arena-Hard methodology would cost around 2M\$, see B.3 for details.. Cost is therefore a strong limitation for the tuning of judges limiting the comparison and tuning of judges to simple aspects such as the choice of the underlying LLM model but excluding other key factors such as the prompt strategy, the output format or the temperature.
In this work, we propose a method to search for optimal judge configurations, including the best prompt parameterization. While prompt tuning approaches such as (Fernando et al., 2023) are related, they do not specifically address tuning judges. A key distinction is that we focus not only on optimizing prompts but also on other judge hyperparameters, such as model selection and temperature. Similarly, Doddapaneni et al. (2024) analyzed the effectiveness of five different prompt strategies for LLM judges. However, their work primarily introduces a benchmark, whereas our approach is centered on systematically tuning and analyzing judges within a search space that includes 80 prompting strategies and a total of 4,480 judge configurations.
We begin by providing background on LLM judges and examining the extent to which improvements can be achieved through scaling alone. This analysis underscores the importance of human agreement as a more efficient metric for differentiating between judges. We then demonstrate how judges can be effectively tuned by optimizing their prompts and other hyperparameters, such as model selection and temperature.
@@ 3.1 Judge
We denote a LLM model as a function $\pi:p\to o$ that produces an output string $o$ given a prompt $p$. A judge compares two models $\pi_{0}$ and $\pi_{1}$ and outputs a score $\Phi(p,\pi_{0}(p),\pi_{1}(p))\in[0,1]$ which is close to zero if $\pi_{0}(p)$ is better than $\pi_{1}(p)$ or close to one otherwise.
A common choice is to use a fixed baseline $\pi^{*}$ for one of the models, typically a frontier model which allows to obtain a score for a model to be evaluated: $\Phi(p,\pi(p))=\Phi(p,\pi^{*}(p),\pi(p))$. Given that the order can sometimes influence the judge decision (Li et al., 2023), previous works have proposed to either randomize the baseline position (Li et al., 2023) or compute the judgement given both orders and average out the result (Li et al., 2024).
@@
One approach to see how well a judge performs is to check how close are its rankings on a grid of models and instructions compared to rankings obtained through human annotations.
Assume we are considering a finite set of models $\mathcal{M}$ and instructions $\mathcal{P}$ and that we are given a list of golden scores for the models denoted $s^{h}\in\mathbb{R}^{\mathcal{M}}$, for instance, the elo-ratings from the Chatbot Arena obtained from human ratings.
To obtain scores from the judge, we average for each model the preference against a baseline $\pi^{*}(p)$, e.g.
$s^{\Phi}_{i}=\mathbb{E}_{p\sim\mathcal{P}}\left[\Phi(p,\pi^{*}(p),\pi_{i}(p))\right].$
One can then use the Spearman correlation between the scores estimated with the judge and the golden score, e.g. $\rho(s^{h},s^{\Phi})=\frac{\text{cov}\left[\text{R}(s^{h}),\text{R}(s^{\Phi})\right]}{\sigma(\text{R}(s^{h}))\sigma(\text{R}(s^{\Phi}))}\in[-1,1]$ where $\text{R}(x)$ and $\sigma(x)$ denotes respectively the rank operator and the standard deviation. Using Spearman correlation is beneficial because it accounts for differences in scale between the golden scores and the judge annotations, focusing on the ranking consistency rather than absolute values. Other metrics which can be used to compare the order of models include Brier score and calibration described (Li et al., 2024).
@@
Another approach to evaluating judges is to measure their agreement with human-annotated preferences in a list of pairwise model battles. Each battle consists of a prompt, a pair of model outputs, and a binary label indicating which output the human annotator preferred.
@@
$\mathbb{E}_{i\sim N}[\Phi(p_{i},o_{i},o^{\prime}_{i})=\Phi^{\text{h}}(p_{i},o_{i},o^{\prime}_{i}))]$ (2)
inline](Verga et al., 2024) also compared with CB Arena, looking at a few prompt variations, no tuning.
@@ 4 Scaling judges
Figure 1: Effect of scaling the LLM judge and increasing the number of instructions on Spearman correlation. In contrast to human agreement, neither Alpaca-Eval, Arena-Hard, nor their union distinguishes the quality difference between 32B and 72B models.
Rather than tuning judges a natural question is: can we just scale them up? LLM judges can be scaled in multiple ways: by increasing the number of parameters of the LLM judge or by using more instructions. Here we investigate how well the Spearman correlation and human agreement metrics scale with both dimensions.
@@
In Fig. 1, we show the scaling behavior when increasing the LLM judge size and the number of instructions. We use Qwen2.5 with a default prompt and compute Spearman correlation on a common set of 26 models available on both Alpaca-Eval and Arena-Hard.
As expected, the judge performance improves when scaling the LLM model and when using more instructions as it allows to cover more areas of the models to evaluate. Alpaca-Eval contains easier prompts and therefore gives better performance to smaller judges compared to Arena-Hard. In contrast, Arena-Hard contains harder and technical questions. This allows to achieve better Spearman correlation as the instructions allow better separability among advanced models, provided that the judge base model is strong enough to measure the performance on this more complex set of instructions.
@@
In Fig. 2, we study the effect of scaling the LLM judge and the number of battles this time on human agreement. Here, we use the same prompt and base LLM models as in the previous paragraph but compute human agreement on the LMSys dataset (lin Chiang et al., 2024).
@@
We have two metrics to evaluate the judge quality: the Spearman correlation and the human agreement. As discussed in the previous paragraph, they behave differently as human agreement is an average of an atomic property (does the judge and human agrees on instruction on average) whereas Spearman correlation is a global metric requiring evaluating many models to compare rankings. Both metrics are correlated, we can see for instance that the ranking w.r.t. the base model size used for the judge is mostly consistent across both Spearman correlation and human agreement in Figs. 1 and 2. For both metrics, we observe that, for this prompt and model family, LLM judges bellow 7B are not able to outperform a simple length baseline which favors the output with the longest answer.
#$\#_{\text{params}}$ Sp. corr. ($\uparrow$) CV($\downarrow$) Hum. agr. ($\uparrow$) CV($\downarrow$)
0.5B 0.09 $\pm$ 0.189 207.60 0.36 $\pm$ 0.006 1.70
1.5B 0.33 $\pm$ 0.137 41.26 0.37 $\pm$ 0.006 1.61
3B 0.82 $\pm$ 0.066 8.11 0.42 $\pm$ 0.006 1.45
7B 0.83 $\pm$ 0.082 9.84 0.45 $\pm$ 0.006 1.33
32B 0.90 $\pm$ 0.052 5.75 0.48 $\pm$ 0.006 1.29
72B 0.88 $\pm$ 0.074 8.43 0.50 $\pm$ 0.006 1.27
Table 1: Comparison of the variability of Spearman-correlation and Human-agreement metrics when using 6500 random annotations and the same default prompt for all model sizes.
In Table 1, we compare the variability of both metrics when using a random set of 6500 instructions [^2] In the case of Spearman correlation, we sample 250 instructions from both Alpaca-Eval and Arena-Hard datasets (which yields to 6500 annotations given that we have 26 models). For LMSys, we just sample 6500 random instructions.. We report the standard deviation and Coefficient of Variation (CV) which is computed as $\sigma/\mu*100$ and indicates the percentage of variation of a metric. It can be seen that while both metrics are correlated, human-agreement allows to differentiate larger models much more easily as the signal to noise ratio of the metric is better. In particular, it allows to statistically distinguish 32B and 72B models whereas those models are tied for Spearman correlation given the number of instructions considered.
In what follows, we therefore use human agreement as the metric to optimize as it allows to rank judge configurations much more cheaply than Spearman correlation and allows to separate judge configurations with much fewer instructions.